OpenStudy (anonymous):

The number of passengers using a railway fell from 216,396 to 197,636 during a 6-year period. Find the annual percentage decrease over this period. Carry out all calculations exactly, but round your final answer to one decimal place. The annual percentage decrease over this period is ? %

OpenStudy (anonymous):

no

OpenStudy (mertsj):

What is it?

OpenStudy (anonymous):

idk but it says its wrong

OpenStudy (mertsj):

216396-197636=18760 18760/216396=.08669 .08669/6=.01445=1.5% per year.

OpenStudy (anonymous):

i need an exact answer. It's for a quiz it should be 1.4% right?